# The time scale of redundancy between prosody and linguistic context

### v2

@@ Abstract
In spoken language, speakers transmit information not only using words, but also via a rich array of non-verbal signals, which include prosody—the auditory features of speech. However, previous studies have shown that prosodic features exhibit significant redundancy with both past and future words. Here, we examine the time scale of this relationship: How many words in the past (or future) contribute to predicting prosody? We find that this scale differs for past and future words. Prosody’s redundancy with past words extends across approximately 3-8 words, whereas redundancy with future words is limited to just 1-2 words. These findings indicate that the prosody-future relationship reflects local word dependencies or short-scale processes such as next word prediction, while the prosody-past relationship unfolds over a longer time scale. The latter suggests that prosody serves to emphasize earlier information that may be challenging for listeners to process given limited cognitive resources in real-time communication. Our results highlight the role of prosody in shaping efficient communication.
@@ 1 Introduction
Auditory features of speech such as pitch, loudness, and tempo—collectively termed prosody—play a crucial role in conveying meaning. Prosody influences sentence-level interpretation, encoding both linguistic and para-linguistic cues relevant to the communicative context Cole (2015); Wagner and Watson (2010); Breen et al. (2010). For example, prosody can signal phrase boundaries, emphasize key elements, transform statements into questions, and express sarcasm, excitement, or doubt. However, some of the cues transmitted by prosody are redundant with the information encoded in the words themselves Wolf et al. (2023). [^1] In both Wolf et al. (2023) and our study, text is used as a proxy to measure information present in the words themselves, also referred to as “segmental information” in the phonology literature. We thus use these terms interchangeably. We also use the term ‘linguistic context’ to describe a word together with the words around it. This raises the question: Why is prosody used if its information is recoverable from the words themselves?
@@
Figure 1: The redundancy between prosody $\mathrm{P}_{t}$and linguistic context $\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}_{\overset{n,m}{\leftrightarrow}}}$—quantified as their mutual information—as a function of the number of words contained in the past ($n$) and/or future ($m$) linguistic context. The values are averaged across 6 prosodic features (see Fig. 3 for the values of each of these features separately).
One possibility is that prosody carries information that is recoverable from long-term past context, but not from short-term past context (see Hypothesis 1 below). Such long-term linguistic information may be challenging for listeners to maintain during real-time communication due to limited human cognitive capacities (e.g., working memory; Gibson, 1998; Lewis et al., 2006; Futrell et al., 2020) and prosody may therefore be used by speakers from an audience-design perspective222‘Audience design’ is sometimes termed ‘listener-oriented’ or ‘intelligibility-oriented’ pressures. Clark and Murphy (1982). For instance, it might be possible to infer which word is most important in the current sentence given long-term linguistic context, but speakers still choose to emphasize the most important word using prosody to help the listener access this information more easily. This explanation implies that the information conveyed by prosody, although redundant with long-term past linguistic information, is locally unique.
Apart from redundancy with past context, however, prosody is also redundant with future words Wolf et al. (2023). Since at any point in time a listener only has access to past context, if prosody is primarily shaped by audience-design, then its redundancy with the future likely reflects its role in helping listeners prepare for, or predict, upcoming words. Assuming prediction mechanisms are mostly tuned for local, incremental comprehension, [^3] This is suggested by recent work on surprisal theory, which shows that limiting the context of language models used to estimate surprisals improves predictive power over reading times (Kuribayashi et al., 2022). we hypothesize that the redundancy between prosody and future context has a shorter scale than with past context (see Hypothesis 2 below).
To summarize our predictions, we thus formulate two main hypotheses:
@@
Long-scale past redundancy. The redundancy between prosody and past linguistic context unfolds across a time scale that is longer than two words.
@@
Short-scale future redundancy. The redundancy between prosody and future linguistic context unfolds across a time scale that is short relative to the past.
To test these hypotheses, we build on the approach of Wolf et al. (2023), quantifying the redundancy between prosody and linguistic context as mutual information, which we estimate using pre-trained language models and a dataset of people reading aloud English audiobooks. We extend their approach to investigate the time scale of this redundancy by limiting the past and future linguistic context available to our models. We vary context lengths parametrically, from 1 to 10 words for both past and future contexts, and analyze how the redundancy changes across time for several prosodic features: pitch, loudness, duration, pause, and prominence. Our main results, averaged across these features, can be seen in Fig. 1. These results confirm that prosody is only redundant with short-term future context (up to 2 words), but with relatively longer-term past context (up to 8 words)—in agreement with our hypotheses. This advances our understanding of the role of prosody in natural spoken communication.
@@ 2 Prosodic Features
Prosodic information can be conveyed via multiple acoustic features of speech, which typically manifest a high degree of co-variation. We next present the features we study here. We independently examine the scale of redundancy with linguistic context for each of these prosodic features.
@@
Pitch is the perceptual dimension over which listeners can order sounds on a scale from low to high. The acoustic correlate giving rise to this perception is the periodicity of sound signals; pitch is thus often measured as the fundamental frequency ($f_{0}$) of the sound. Pitch is (arguably) the hallmark feature of prosody, with its contours, having been extensively studied and characterized (Pierrehumbert, 1980; Silverman et al., 1992; Jun, 2006). These contours carry contextual information that can signal a wealth of information, including, in stress-accent languages like English, emphasizing specific words, signaling boundaries, speech act type, and speaker’s intent (like interrogation, sarcasm, and the affective state of the speaker). Some typical pitch curves are the rise of pitch towards the last word of a question (e.g. in yes/no questions in American English), rise and then fall of pitch on a specific word to emphasize it, and a fall toward the end of phrases.
@@
Loudness is the perceptual dimension over which listeners can order sounds on a scale from quiet to loud. The acoustic correlate giving rise to this perception is sound pressure, being measured as the intensity of acoustic energy. The loudness of speech can be used to transmit information such as emphasize important words or convey emotion. The correlation between pitch and loudness is partly explained by vocal production constraints; producing speech with higher energy is helpful for raising and better controlling the fundamental frequency. However, loudness variations may also convey independent information from pitch.
@@
A word’s duration is the difference between its offset (end) and onset (start) times. The relationship between word duration and linguistic information has long been studied as a signature of efficiency in communication such that more predictable words are reduced to a shorter duration (Jurafsky et al., 1998; Bell et al., 2009; Seyfarth, 2014; Coupé et al., 2019; Pimentel et al., 2021). Further, elongating a word is a common way to emphasize it, or signal prosodic boundary. Duration is thus also highly correlated with pitch and loudness in natural speech, but it can also be used independently to convey meaning, or to compress words of low information content.
@@
A word’s pause is the time difference between its offset (end) time and the next word’s onset (start), being another way to emphasize an important word in context, or to signal phrase boundaries (Hawkins, 1971). In contrast to phrase boundaries, within the phrase speech tends to be ’connected’ such that there are usually no pauses between words; i.e., most pauses are of zero seconds.
@@
Prosodic prominence is a term that describes how salient a linguistic entity—in our case, a single word—is perceived relative to the words surrounding it in an utterance Terken and Hermes (2000). Unlike the previously described prosodic features, prominence is a higher-level percept in the sense that it is not elicited by a single acoustic dimension. The perception of prominence is affected by multiple acoustic features Cole et al. (2010)—elongating a word’s duration, increasing the speech energy or modulating the $f_{0}$ contour of a specific word can all make this word be perceived as more prominent in context. Although other acoustic features, like timbre, can also affect prominence, and factors like word frequency influence it as well Cole et al. (2010), a combination of duration, loudness and pitch has been proposed as an effective acoustic measure to quantify prosodic prominence Talman et al. (2019), which we use here.
@@ 3 Redundancy between Prosody and Linguistic Context
This paper concerns the time scale of the redundancy between prosody and linguistic context, where ‘linguistic context’ here refers to the segmental information of an utterance, represented in our experiments as text. In this section, we first explain how this redundancy can be formalized as a mutual information, following Wolf et al. (2023). We then expand on this framework by discussing how context-length manipulations allow us to investigate the time scale aspect of this redundancy.
@@ 3.1 Redundancy as Mutual Information
Let ${{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}}$ be a prosody-valued random variable, which takes values ${{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{p}_{t}}}\in\mathbb{R}$. Further, let ${{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}$ be a words-valued random variable, which takes values ${{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}\in\mathcal{W}^{*}$, where $\mathcal{W}$ is a language’s lexicon. We follow Wolf et al. (2023) in formalizing the redundancy between prosody and linguistic context as the mutual information: ${\mathrm{MI}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}};{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}})$. Under a few technical assumptions (e.g., the good mixed-pair assumption, see Wolf et al., 2023 for details), we can write this value as:
@@
In this equation, unconditional entropy $\mathrm{H}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}})$ serves as a baseline, representing how much uncertainty there is about ${{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}}$. In turn, conditional entropy $\mathrm{H}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}}\mid{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}})$ represents how much uncertainty remains about ${{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}}$ once we know the context ${{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}$. Their difference then represents how much information ${{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}$ contains about ${{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}}$ (and vice versa).

### v3

@@ Abstract
In spoken communication, information is transmitted not only via words, but also through a rich array of non-verbal signals, including prosody—the non-segmental auditory features of speech. Do these different communication channels carry distinct information? Prior work has shown that the information carried by prosodic features is substantially redundant with that carried by the surrounding words. Here, we systematically examine the time scale of this relationship, studying how it varies with the length of past and future contexts. We find that a word’s prosodic features require an extended past context (3-8 words across different features) to be reliably predicted. Given that long-scale contextual information decays in memory, prosody may facilitate communication by adding information that is locally unique. We also find that a word’s prosodic features show some redundancy with future words, but only with a short scale of 1-2 words, consistent with reports of incremental short-term planning in language production. Thus, prosody may facilitate communication by helping listeners predict upcoming material. In tandem, our results highlight potentially distinct roles that prosody plays in facilitating integration of words into past contexts and in helping predict upcoming words.
@@ 1 Introduction
Auditory features of speech such as pitch, loudness, and tempo—collectively termed prosody—play a crucial role in conveying meaning. Prosody influences sentence-level interpretation, encoding both linguistic and para-linguistic cues relevant to the communicative context Cole (2015); Wagner and Watson (2010); Breen et al. (2010). For example, prosody can signal phrase boundaries, emphasize key elements, transform statements into questions, and express sarcasm, excitement, or doubt. However, much of the information conveyed by prosody is redundant with the information encoded in the words themselves Wolf et al. (2023)—that is, it is possible to predict a word’s prosodic features from its linguistic context, i.e., its surrounding words. [^1] In both Wolf et al. (2023) and our study, text is used as a proxy to measure information present in the words themselves, which we refer to as ”linguistic information”(also termed “segmental information” in the phonology literature).
@@
Figure 1: Redundancy ($\mathrm{MI}$) between prosody ($\mathrm{P}_{t}$) and linguistic context ($\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}_{\overset{n,m}{\leftrightarrow}}}$) as a function of the number of words contained in the past ($n$) and/or future ($m$) context. The values are averaged across 6 prosodic features (see Fig. 3 for each of these features separately). The first column is used to evaluate Hypothesis 1; the first row, Hypothesis 2 (see section § 1). Time scales are defined as the time after which the $\mathrm{MI}$ does not increase.
Past work, however, only quantified the redundancy between prosody and the entire linguistic context. The question of the dynamics of this redundancy is thus left open. Here, we investigate the time scale of redundancy between prosodic features of a given word and its linguistic context by systematically varying the length of past or future context from 0 to 9 words in each direction (0 includes only the word whose prosody is being tested).
A priori, what might one expect regarding the time scale of redundancy between prosodic features and past linguistic context? During language comprehension, context is often critical to interpreting the current word and integrating it into the linguistic representations built thus far (Hale, 2001; Levy, 2008). However, human memory is limited: contextual information decays over time, and connecting an incoming word to one far back is costly Gibson (1998); Lewis et al. (2006); Futrell et al. (2020). Given these memory limitations, prosody may be critically helpful in conveying information that is redundant with past context but is locally unique, such that if a listener’s representation of the past context is imperfect, they would benefit from the prosodic information contained by the word itself. We therefore hypothesize that prosody carries information that is recoverable from long-term past context, but not from the short-term past (see Hypothesis 1 below).
With respect to future linguistic context, lexical and syntactic planning in language production is known to be relatively restricted Brown-Schmidt and Konopka (2008); Bock et al. (1994): we may have a general idea of what we want to say next, but the particular words and constructions are selected incrementally, often with revisions. As a result, how we pronounce a current word and what prosodic features we associate with it should not be affected by long-term future linguistic context. We therefore hypothesize that prosody carries information that is not recoverable from long-term future context (see Hypothesis 2 below).
The above logic is laid out in terms of “short” vs. “long” contexts, but what constitutes a short or long time scale during language processing? Following recent work in neuroscience Jain and Huth (2018); Regev, Casto, et al. (2024); Shain, Kean, et al. (2024) as well as work investigating intonational units in prosody itself (Inbar et al., 2020), we set the cutoff between “short” and “long” at roughly 1-2 words, although see section § 6.1 for more discussion.
We thus formulate two main hypotheses:
@@
Long-scale past redundancy. The redundancy between prosody and past linguistic context unfolds across a long time scale (longer than 1-2 words).
@@
Short-scale future redundancy. The redundancy between prosody and future linguistic context unfolds across a short time scale (of 1-2 words).
To test these hypotheses, we quantify the redundancy between prosody and linguistic context as their mutual information, which we estimate using pre-trained language models, following Wolf et al. (2023). Notably, we extend their approach to investigate the time scale of this redundancy by jointly varying context lengths parametrically, from 0 to 9 words for both past and future contexts. [^2] In addition to examining the effects of the length of the past and future contexts separately, we also considered different combinations of past and future context lengths in order to identify the most predictive combination. We then analyze how redundancy changes across time for several commonly discussed prosodic features: pitch, loudness, duration, pause, and prominence.
Our main results, averaged across all prosodic features considered are shown in Fig. 1 (see Fig. 3 for feature-specific results). In line with our hypotheses, these results demonstrate that prosody is redundant with relatively long-term past linguistic context (up to 8 words), but is only redundant with short-term future context (up to 2 words). This finding advances our understanding of the dynamics of the redundancy between linguistic context and prosody, with potential implications for the role of prosody in natural spoken communication.
@@ 2 Prosodic Features
Prosodic information is conveyed through multiple acoustic features of speech. Here, we examine the redundancy between linguistic context and: pitch, loudness, duration, pause and two versions of prominence (as detailed in section § 4.1). We chose these prosodic features as they have been commonly investigated in prior prosody research (e.g., Breen et al., 2010; Gibson, 1998; Cole, 2015).
@@
Pitch is the perceptual dimension over which listeners order sounds on a scale from low to high. The acoustic correlate giving rise to this perception is the periodicity of sound signals; pitch is thus often measured as the fundamental frequency ($f_{0}$) of the sound. Pitch is (arguably) the hallmark feature of prosody, having been extensively studied and characterized (Pierrehumbert, 1980; Silverman et al., 1992; Jun, 2006). In stress-accent languages like English, pitch contours carry contextual information that can signal a wealth of information, including emphasizing specific words, signaling boundaries, speech act type, and the speaker’s intent (e.g., interrogation, sarcasm, or affective state). Some typical pitch curves are the rise of pitch towards the last word of a question in yes/no questions in American English, the rise and then fall of pitch on a specific word to emphasize it, and a fall toward the end of phrases. Note that we do not study here the relationship between pitch and lexical identity, as in tonal-languages. However, see Wilcox et al. (2025), who use similar methods to study typological pitch variation across languages.
@@
Loudness is the perceptual dimension over which listeners can order sounds on a scale from quiet to loud. The acoustic correlate giving rise to this perception is sound pressure, being measured as the intensity of acoustic energy. The loudness of speech can be used to transmit information, such as to emphasize important words or convey emotion. The correlation between pitch and loudness is partly explained by vocal production constraints; producing speech with higher energy helps raise and better control the fundamental frequency. However, loudness variations may also convey independent information from pitch.
@@
A word’s duration is the difference between its offset (end) and onset (start) times. The relationship between word duration and linguistic information has long been studied as a signature of efficiency in communication, such that more predictable words are reduced to a shorter duration (Jurafsky et al., 1998; Bell et al., 2009; Seyfarth, 2014; Coupé, Oh, et al., 2019; Pimentel et al., 2021). Further, elongating a word is a common way to emphasize it, or signal prosodic boundary. Duration is thus also highly correlated with pitch and loudness in natural speech, but it can also be used independently to convey meaning, or to compress words of low information content.
@@
A pause following a word is the time difference between the word’s offset (end) time and the next word’s onset (start), being another way to emphasize an important word in context, or to signal phrase boundaries (Hawkins, 1971). In contrast to phrase boundaries, within the phrase speech tends to be ‘connected’ such that there are usually no pauses between words; i.e., most pauses are of zero seconds (see section § 4.1).
@@
Prosodic prominence is a term that describes how salient a linguistic entity—in our case, a single word—is perceived relative to the words surrounding it in an utterance Terken and Hermes (2000). Unlike the previously described prosodic features, prominence is a higher-level percept in the sense that it is not elicited by a single acoustic dimension. The perception of prominence is affected by multiple acoustic features Cole et al. (2010)—elongating a word, increasing the speech energy or modulating the $f_{0}$ contour of the word can all make this word be perceived as more prominent in context. Although other acoustic features, like timbre, can also affect prominence, and factors like word frequency influence it as well Cole et al. (2010), a combination of duration, loudness and pitch has been proposed as an effective acoustic measure to quantify prosodic prominence Talman et al. (2019), which we use here.
Notably, the prosodic features above present a high degree of correlation in natural speech Ladd (1996). For instance, as noted above, producing a higher pitch generally requires greater vocal effort, resulting in increased intensity; similarly, it may take longer to reach a higher pitch target, creating dependencies between pitch, loudness, and duration. Furthermore, prominence is strongly correlated to pitch, loudness and duration, as it is directly computed using a combination of these acoustic measures. We did not model these correlations explicitly; instead, we quantified the redundancy of each feature with linguistic context separately. This design choice does not affect our central research question, which concerns the time scale of these redundancies. Furthermore, although these prosodic features are correlated, the differences observed across them in our results may indicate that they also carry distinct information.
@@ 3 Redundancy between Prosody and Linguistic Context
This paper concerns the time scale of the redundancy between prosody and linguistic context, where ‘linguistic context’ here refers to the segmental information of an utterance, represented in our experiments as text. In this section, we first explain how this redundancy can be formalized as a mutual information, following Wolf et al. (2023). We then expand on this framework by discussing how context-length manipulations allow us to investigate the time scale of this redundancy.
@@ 3.1 Redundancy as Mutual Information
Let ${{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}}$ be a prosody-valued random variable, which takes values ${{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{p}_{t}}}\in\mathbb{R}$. Further, let ${{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}$ be a words-valued random variable, which takes values ${{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}\!\in\!\mathcal{W}^{*}\!\!$, where $\mathcal{W}$ is a language’s lexicon. We follow Wolf et al. (2023) in formalizing the redundancy between prosody and linguistic context as the mutual information: ${\mathrm{MI}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}};{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}})$. Under a few technical assumptions (e.g., the good mixed-pair assumption, see Wolf et al. for details), we can write this as:
@@
In this equation, unconditional entropy $\mathrm{H}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}})$ serves as a baseline, representing how much uncertainty there is about ${{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}}$, in general. In turn, conditional entropy $\mathrm{H}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}}\mid{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}})$ represents how much uncertainty remains about ${{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}}$ once we know ${{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}$. Their difference then represents how much information ${{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}$ contains about ${{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}}$ (and vice versa).